OpenStudy (anonymous):

select the ratios that from a proportion. 3:6 and 5:10 4/12 and 6/18 20/5 and 10/3 24:36 and 8:9

OpenStudy (demonchild99):

so like what do you have to do?....find one that the numbers go in equaly?

OpenStudy (igreen):

3 / 6 = ? 5 / 10 = ? @noterdame13

OpenStudy (igreen):

@demonchild99 It's asking which two proportions are equal.

OpenStudy (igreen):

To find out we can divide them and convert them into decimal form.

OpenStudy (demonchild99):

ohhh ok....lol

OpenStudy (anonymous):

so what are the answers

OpenStudy (midhun.madhu1987):

Equal ratios are called Proportion.

OpenStudy (igreen):

Divide these. @noterdame13 3 / 6 = ? 5 / 10 = ?

OpenStudy (anonymous):

there both 2

OpenStudy (demonchild99):

we are not allowed to give direct answer @noterdame13 ...you must find it yourselve we can help but no direct answer

OpenStudy (igreen):

No, you did 6 / 3 and 10 / 5, it's 3 / 6 and 5 / 10.

OpenStudy (anonymous):

there not equal

OpenStudy (demonchild99):

im going to wing it....do you see any ove the equations above that are divided equally

OpenStudy (anonymous):

22222@@2222@@@@@@@@@@iGreen

OpenStudy (demonchild99):

@iGreen

OpenStudy (demonchild99):

look above there is one answer where it can be divided by the numerator or watever you wnna call it

OpenStudy (demonchild99):

am i doing it right? @iGreen

OpenStudy (igreen):

Yes, you can say they both equal 2..which means 3:6 and 5:10 form a proportion.

OpenStudy (igreen):

Now divide: 4 / 12 and 6 / 18 @noterdame13

OpenStudy (demonchild99):

so there is only one answer that you can go with...ok cuz thats what i was thinking

OpenStudy (anonymous):

demon child why did you dimond clark was a boy i founed out that she is a girl

OpenStudy (igreen):

..............

OpenStudy (anonymous):

there are both 3

OpenStudy (igreen):

Yep, so 4:12 and 6:18 also form a proportion. So your answers are 3:6 and 5:10 and 4:12 and 6:18.

OpenStudy (igreen):

What are you guys talking about..?

OpenStudy (igreen):

@noterdame13 You have your answers :) A & B